Suppose we create a regression model where height (in feet) is used to predict weight (in lbs). Then, we realize that, as scientists, we should use the metric system. The heights are converted to meters and the weights to kilograms, then we redo the analysis. What is the effect on the slope in regression model? (A) the sign of the slope will change (B) the magnitude of the slope will change (C) the sign and magnitude of the slope will both change (D) there will be no effect on the slope
Answer #1

Weight = c + k * height

1 lbs = 0.454 Kg

1 foot = 0.305 meters

0.454 * weight = c + k * height * 0.305

New slope will be = k * 0.305 / 0.454 = 0.67 * k

Option-B) the magnitude of the slope will change is the correct answer.